HUBBLE SPACE TELESCOPE DAILY REPORT #2500 PERIOD COVERED: 0000Z (UTC) 11/12/99 - 0000Z (UTC) 11/15/99 Daily Status Report as of 319/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of STIS/CCD 8545 (POMS Test Proposal: STIS non-scripted parallel proposal continuation VI) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. As described IN HSTAR 7355 and 2.1, re-acquisitions for the third and fourth iterations of this proposal failed and the take data flag remained down. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.2 Completed Four Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. As described IN HSTAR 7355 and 2.1, re-acquisitions for the third and fourth iterations of this proposal failed and the take data flag remained down. Otherwise, the observations were executed as scheduled, and no further anomalies were reported. 1.3 Completed Four Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described IN HSTAR 7355 and 2.1, re-acquisitions for the third and fourth iterations of this proposal failed and the take data flag remained down. Otherwise, the observations were completed as planned, and no further problems were reported. 1.4 Completed Two Sets of STIS/CCD 8409 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. As described IN HSTAR 7355 and 2.1, re-acquisitions for the second iteration of this proposal failed and the take data flag remained down. Otherwise, the observations completed with no further problems. 1.5 Completed STIS/CCD/MA1/MA2 8112 (Probing the Galactic Halo and Beyond with Young Supernovae) The Space Telescope Imaging Spectroscope (CCD/MA1/MA2) was used for a target of opportunity program to obtain STIS echelle spectra of a bright new supernova {V<14} to characterize the ionization state, gas-phase abundances, depletion on dust grains, metallicity, and gas kinematics in the Galactic halo, the ISM and halo of the SN's host galaxy, and, if the properties of the sightline are favorable, in the intervening intergalactic medium. The observations completed nominally. 1.6 Completed WF/PC-2 8272 (The Horizontal Branches of the M31 Dwarf Spheroidal Companions And V & VI) The WF/PC-2 was used to make observations of stars in the horizontal branches of the M-31 dwarf spheroidal companions galaxies Andromeda V & VI. These dwarf spheroidal galactic companions show a surprising diversity in the epochs of their major star formation episodes, extending from ~15 Gyr to ~3 Gyr. The observations were executed as scheduled, and no anomalies were reported. 1.7 Completed Four Sets of STIS/CCD 8549 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations as part of a non-scripted POMS Test Proposal. As described IN HSTAR 7355 and 2.1, re-acquisitions for the second, third, and fourth iterations of this proposal failed and the take data flag remained down. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.8 Completed STIS/MA2 8429 (Cycle 8 Plate Scale Verification) The Space Telescope Imaging Spectrograph (MA2) was used to conduct a Cycle-8 plate scale calibration verification. The calibration was completed as scheduled, and no anomalies were reported. The following scheduled proposals were not executed due to the Zero Gyro Sun Point entry described in 2.3: Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) Three Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) STIS/CCD 8409 (CCD Bias Monitor C8) Nine Sets of STIS/CCD 8549 (POMS Test Proposal: STIS Non- scripted Parallel Proposal Continuation III) STIS/CCD/MA1 8108 (Search for the Origin of Supersonic Turbulence Observed in the Upper Equatorial Atmosphere of JupiterP WF/PC-2 8326 (Colliding Stellar Winds and Proto-Planetary Dynamics) WF/PC-2 8114 (Fundamental Problems in Plasma Astrophysics) Two Sets of WF/PC-2 8338 (The Evolution of the Host Galaxies of Powerful Radio Sources) STIS/MA2 7752 (The Distribution and Evolution of Lyman-Alpha Forest Cloud Sizes) WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) WF/PC-2 8290 (Weather, Moons, and Orbit of the Brown Dwarf Gl 229B) WF/PC-2 8175 (A Survey of Gravitational Lenses as Cosmological Tools) STIS/MA1 8428 (STIS FUV-MAMA Cycle 8 Flats) WF/PC-2 8398 (Saturn's Rings) WF/PC-2 8146 (A Search for Binary L-Dwarfs) WF/PC-2 8194 (Evolution of High-Redshift Seyfert Galaxies) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 2 Successful: 2 Scheduled Re-acquisitions: 11 Successful: 7 Per HSTAR, the acquisition at 317/000543Z acquired fine lock, but lost lock at 317/001251Z. The vehicle lost and re-acquired fine lock eleven times between 317/001251Z and 317/003951Z. FHST maps at 317/0040Z and at 317/0218Z showed high attitude errors. The re-acquisition at 317/0143Z failed. An ARU/PRT was uplinked at 317/0316Z, but the following re-acquisition at 317/0320Z also failed. Other maps continued to show high attitude errors. Other ARU/PRTs were uplinked, but the re-acquisitions at 317/0455Z and 317/0945Z also failed. Proposals list in 1.1, 1.2, 1.3, 1.4, and 1.7 were affected. 2.2 FHST Updates: Scheduled: 6 Successful: 6 2.3 Operations Notes: The SSR EDAC errors were cleared six times per ROP SR-1. STIS MCE-2 reset at 317/015853Z while the low voltage was on and while outside any SAA interval. Using ROP NS-12, the STIS flight software error counter was cleared at 317/0207Z. MAMA-2 was recovered via normal SMS commanding at 317/1119Z. According to HSTAR 7357, Gyro 1 experienced an anomaly which resulted in HST entry into Zero Gyro Sun Point (ZGSP) at 317/132939Z. The gyro anomaly manifested itself as an exponential motor current increase from 160 mA to 367 mA over a 15 minute period. After peaking at 367 mA at 317/1334Z, the motor current decreased to a steady-state of 347 mA. Gyro 1 analog rate showed a step increase from 1 deg/hr to 65 deg/hr at the time of ZGSP entry. Gyro 1 was no longer responsive after this time, although an attempt was made at 317/180740Z to recover with a power cycle. Gyro 1 was kept on long enough to collect duty cycle data to support the anomaly analysis. This investigation continues. HST remains in ZGSP with nominal performance, except for the Battery Rate of Charge problem mentioned in the next paragraph. An operations note was generated which provide guidelines for monitoring HST in ZGSP. The Battery Rate of Charge error counter (QROCERC) failed (flagged high) at 318/190821Z, after having entered orbit day at 318/1858Z. HSTAR 7358 was written. Because the Rate of Charge Test failed, a safing macro initiated load shedding, which safed the NSSC-1 and powered off the FGEs and FHSTs. The FHSTs were powered on at 319/0253Z; the FGEs, at 319/0359Z. The Battery Rate of Charge Safemode Test was re-enabled at 319/0823Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW